# Distorting Embedding Space for Safety: A Defense Mechanism for Adversarially Robust Diffusion Models

### v1

@@ Abstract
Text-to-image diffusion models show remarkable generation performance following text prompts, but risk generating Not Safe For Work (NSFW) contents from unsafe prompts. Existing approaches, such as prompt filtering or concept unlearning, fail to defend against adversarial attacks while maintaining benign image quality. In this paper, we propose a novel approach called Distorting Embedding Space (DES), a text encoder-based defense mechanism that effectively tackles these issues through innovative embedding space control. DES transforms unsafe embeddings, extracted from a text encoder using unsafe prompts, toward carefully calculated safe embedding regions to prevent unsafe contents generation, while reproducing the original safe embeddings. DES also neutralizes the nudity embedding, extracted using prompt “nudity”, by aligning it with neutral embedding to enhance robustness against adversarial attacks. These methods ensure both robust defense and high-quality image generation. Additionally, DES can be adopted in a plug-and-play manner and requires zero inference overhead, facilitating its deployment. Extensive experiments on diverse attack types, including black-box and white-box scenarios, demonstrate DES’s state-of-the-art performance in both defense capability and benign image generation quality. Our model is available at https://github.com/aei13/DES.
@@ 1 Introduction
Recent advances in diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020), including Stable Diffusion (Rombach et al., 2022) and DALL-E (Betker et al., 2023), have demonstrated remarkable capabilities in various image generation tasks such as text-to-image (T2I) synthesis and image inpainting (Brooks et al., 2023). However, these models exhibit vulnerability to malicious exploitation, specifically in generating harmful or Not Safe For Work (NSFW) content (Yang et al., 2024c). While these models employ safety checkers or prompt filtering to prevent NSFW content generation (Rando et al., 2022), attackers can bypass them through adversarial attacks, which are challenging to defend against. The widespread public accessibility of these models raises significant ethical concerns regarding their responsible deployment and utilization (Truong et al., 2024).
Figure 1: Distorting Unsafe Embedding Space. Our method distorts the unsafe embedding space by transforming unsafe embeddings into a safe embedding region. This transformation ensures that even embeddings derived from unsafe or adversarial prompts result in the generation of benign content.
Figure 2: Performance comparison of our DES over other defense mechanisms. DES offers the most accurate and high-quality solution in terms of attack success rate (ASR) and Fréchet Inception Distance (FID) (Heusel et al., 2017), while also being cost-effective in training time. The relative sizes of the circles illustrate training time. ASR values are computed by averaging across representative adversarial attacks, such as SneakyPrompt (Yang et al., 2024c), MMA (Yang et al., 2024a), I2P (Schramowski et al., 2023), Ring-A-Bell (Tsai et al., 2024), and P4D (Chin et al., 2024).
Recent studies have addressed this vulnerability through post-hoc defense and unlearning-based defense. Post-hoc defense methods implement filtering mechanisms at various stages (Rando et al., 2022; Liu et al., 2025). However, the inherent ambiguity of natural language enables attackers to encode unsafe content without explicit NSFW terms, making prompt-based filtering ineffective (Yang et al., 2024c). The vulnerability is particularly severe in open-source models, where attackers can leverage gradient-based optimization to construct adversarial prompts (Yang et al., 2024a). This creates two distinct attack scenarios: black-box attacks against closed-source models and more sophisticated white-box attacks against open-source models, both requiring robust defense mechanisms. Machine unlearning presents an alternative defense strategy by attempting to eliminate unsafe content generation capabilities from diffusion models (Gandikota et al., 2023). While it aims to prevent unsafe content generation, it faces significant limitations as most unlearning methods fail to achieve robust defense against unsafe image generation, leaving the model vulnerable to reactivation through adversarial attacks (Zhang et al., 2025), while also struggling to maintain high-quality benign content generation (Basu et al., 2023). These limitations undermine the effectiveness of unlearning-based approaches.
@@
To address these challenges, we propose a novel approach, Distorting Embedding Space (DES), a defense framework that satisfies both robust protection against NSFW content generation and high-quality safe content generation. Unlike existing methods that struggle with implicit NSFW representations, DES uniquely controls the comprehensive embedding space to capture various implicit NSFW representations, effectively defending against both black-box and white-box attacks. Our framework first transforms unsafe embeddings into a designated safe region. Since this transformation can potentially affect safe embeddings and degrade benign image generation quality, DES simultaneously trains the text encoder to reproduce the original safe embeddings, as illustrated in Figure 1. This dual-objective is achieved by three loss functions: Unsafe Embedding Neutralization (UEN), Safe Embedding Preservation (SEP) with Proximity-Aware Loss Adjustment (PALA), and Nudity Embedding Neutralization (NEN). These loss functions enable DES to successfully generate safe content from unsafe prompts while maintaining high-quality image generation. As demonstrated in Figure 2, DES significantly outperforms existing defense mechanisms in terms of ASR and FID. Furthermore, DES offers remarkable efficiency in both training and inference: it requires only 90 seconds for training, introduces zero inference overhead, and can be deployed in a plug-and-play manner across different T2I models.
@@
Our contributions are as follows: 1. We propose DES, a novel defense framework that uniquely controls the text embedding space through UEN, SEP, PALA, and NEN, achieving state-of-the-art (SOTA) defense performance against NSFW adversarial attacks while maintaining benign image generation quality. 2. We develop a practical, plug-and-play solution that requires efficient training with zero-inference overhead through simple text encoder weight swapping, enabling easy deployment in real-world applications. 3. We provide extensive evaluations in both black-box and white-box attack scenarios and extensive analyses of embedding space distortion, demonstrating DES’s effectiveness and robustness.
@@ 2.1 Adversarial Attacks
Text-to-image (T2I) diffusion models, trained on large-scale uncurated datasets, can generate inappropriate content when given unsafe prompts (Schramowski et al., 2023). While prompt filtering effectively blocks such prompts, recent studies reveal that adversarial attacks can bypass these filters (Pham et al., 2023; Deng & Chen, 2023; Chin et al., 2024). These attacks have become increasingly sophisticated, employing various optimization techniques to circumvent safety filters. SneakyPrompt (Yang et al., 2024c) leverages reinforcement learning to craft adversarial prompts that generate images semantically similar to target prompts, MMA-diffusion (Yang et al., 2024a) employs gradient-based optimization to create prompts that closely resemble target prompts. Ring-A-Bell (Tsai et al., 2024) uses a genetic algorithm to discover malicious prompts similar to combinations of normal embeddings and extracted nudity embedding. These attacks effectively bypass safety filters by exploiting unsafe embedding subspaces inherited from the training data in white-box and black-box scenarios.
Figure 3: Overview of DES framework. During target vector generation phase, DES searches safe-unsafe vector pairs and creates target vectors by subtracting the “nudity” direction from minimum similarity safe vectors. In training phase, DES aligns unsafe vectors with target vectors and maintains safe vectors by aligning both their current and nudity direction-integrated states with the originals. It also aligns the “nudity” vector with the unconditioned vector to neutralize its semantics. Here, $e$ is the original vector from the original text encoder, $\tilde{e}$ denotes the current vector from the training text encoder.
@@ 2.2 Defense Methods
2.2.1 Post-hoc Defense Methods
Several defense mechanisms have been proposed to address these vulnerabilities (OpenAI, 2024; Amazon Web Services, 2024), generally relying on embedding-based contextual analysis (Yang et al., 2024b; Yoon et al., 2024). GuardT2I (Yang et al., 2024b) leverages a Large Language Model for NSFW detection through embedding interpretation, while SAFREE (Yoon et al., 2024) proposes training-free filtering based on distances between masked embeddings and unsafe concepts. However, these methods require additional model training or introduce inference overhead. Furthermore, these approaches struggle to detect unsafe content in ambiguous expressions and remain vulnerable to white-box attacks. In contrast, DES operates directly on the text encoder without requiring additional models or computational overhead, while effectively handling unsafe prompts in both white-box and black-box scenarios through its embedding space control.
2.2.2 Unlearning-based Defense Methods
Recent approaches explore machine unlearning (Gandikota et al., 2024; Zhang et al., 2024a; Lyu et al., 2024). ESD (Gandikota et al., 2023) develops a concept erasure mechanism that steers model outputs away from specific concepts. SalUn (Fan et al., 2023) suggests saliency-based unlearning, which assigns random concepts to specific concepts to unlearn the concept. However, these UNet-based methods remain vulnerable to adversarial attacks (Sharma et al., 2024; Zhang et al., 2025) or compromise image generation quality. AdvUnlearn (Zhang et al., 2024b) attempts to address these issues by optimizing text encoder, incorporating adversarial training. Nevertheless, it suffers from degraded image quality, a common limitation of adversarial training that compromises model performance (Tsipras et al., 2018). In contrast, DES overcomes these limitations through embedding space control rather than UNet modification or adversarial training, achieving robust defense while maintaining generation quality and enabling plug-and-play deployment without model-specific modifications.
@@ 3 Proposed Methods
Figure 3 provides an overview of DES, illustrating the target vector generation and training phases. The first phase calculates the desired transformation targets for unsafe prompts, ensuring they are directed to locations that provide safe meanings without significantly disrupting the safe embeddings. During the training phase, the text encoder is fine-tuned to unlearn unsafe information while explicitly preserving the information of safe embeddings.
@@ 3.1 Target Vector Generation Phase
To prevent unsafe content generation, we propose transforming unsafe embeddings into the safe embedding region or to locations significantly different from their original positions. This phase involves identifying optimal target safe vectors that are most dissimilar to unsafe vectors, as greater dissimilarity is assumed to enhance robustness by increasing embedding space distortion. Therefore, we search through all safe vectors to identify those with minimum cosine similarity to each unsafe vector. An analysis of robustness based on dissimilarity is provided in Appendix G.1. This selection procedure can be formalized as:
$\bar{e}_{s,i}=\arg\min_{e_{s,i}}\left(\frac{e_{u,i}\cdot e_{s,i}}{\|e_{u,i}\|\|e_{s,i}\|}\right),$ (1)
where $i\!=\!1,\ldots,M$ indexes $M$ vectors, and $e_{s,i}$, $e_{u,i}$, and $\bar{e}_{s,i}$ denote safe, unsafe, and selected safe vectors, respectively. Examples of safe prompts corresponding to these selected vectors are provided in Appendix F.
Figure 4: Cosine similarity distributions between the “nudity” vector and other vectors. Selected safe vectors initially exhibit positive similarities, which decrease as the “nudity” direction, scaled by $s_{g}$, is subtracted.
We then observe the similarity between the selected safe vectors and the “nudity” vector. Interestingly, as shown in Figure 4, selected safe vectors have positive correlations with “nudity”. While the selection strategy improves robustness, we propose further enhancement by subtracting the “nudity” direction while using the selected vectors as basis vectors, creating target vectors $\hat{e}_{s,i}$ that are anti-correlated with the unsafe vector. This subtraction step is represented as:
$\hat{e}_{s,i}=\bar{e}_{s,i}-s_{g}\frac{e_{n}}{\|e_{n}\|},$ (2)
where $e_{n}$ is the “nudity” vector and $s_{g}$ is a scaling factor. This ensures that unsafe vectors are directed away from the “nudity” direction. However, excessive subtraction can cause performance degradation if the embeddings deviate too far from the learned embedding space. Therefore, we adopt $s_{g}$ to control the scale of subtraction. The green-hued distributions in Figure 4 demonstrate the successful creation of these anti-correlated vectors, which serve as more effective transformation targets for unsafe vectors. We provide a detailed description of the target vector generation phase in Algorithm 1.
@@
0:  Pretrained text encoder $\mathcal{E}_{\phi_{o}}$, safe prompts $\mathcal{P}_{s}=\{p_{s,1},\ldots,p_{s,M}\}$, unsafe prompts $\mathcal{P}_{u}=\{p_{u,1},\ldots,p_{u,M}\}$, nudity prompt $p_{n}$, scale factor $s_{g}$
1:  $e_{n}\leftarrow\mathcal{E}_{\phi_{o}}(p_{n})$ // Extract nudity vector
2:  $\mathcal{D}\leftarrow\emptyset$
@@
4:     $e_{u,i}\leftarrow\mathcal{E}_{\phi_{o}}(p_{u,i})$ // Extract unsafe vectors
5:     $\bar{e}_{s,i}$ is computed by Eq. (1) // Select safe vectors
6:     $\hat{e}_{s,i}$ is computed by Eq. (2) // Nudity subtraction
7:     $\mathcal{D}\leftarrow\mathcal{D}\cup\{(\hat{e}_{s,i},p_{u,i},p_{s,i})\}$ // Save pairs
@@ 3.2 Training Phase
3.2.1 Unsafe Embedding Neutralization (UEN)
@@
In the text embedding space, unsafe embeddings should not occupy positions associated with unsafe content. They should be transformed into safe embedding regions or moved from their original positions by fine-tuning the text encoder weights. Specifically, we propose the UEN loss, which minimizes the cosine similarity between the current unsafe vectors $\tilde{e}_{u,i}$ and the target safe vectors $\hat{e}_{s,i}$:
$\mathcal{L}_{u}=\frac{1}{B}\sum_{i=1}^{B}\left(1-\frac{\tilde{e}_{u,i}\cdot\hat{e}_{s,i}}{\|\tilde{e}_{u,i}\|\|\hat{e}_{s,i}\|}\right),$ (3)
@@
where $i\!=\!1,\ldots,B$ represents each embedding in a mini-batch, and $B$ denotes the batch size for each iteration. Each of $\tilde{e}_{u,i}$ and $\hat{e}_{s,i}$ represents an $i$-th embedding vector in a mini-batch. It aligns unsafe vectors with target vectors, avoiding their original positions. In particular, unsafe vectors become anti-correlated with the unsafe concept “nudity,” ensuring its removal from unsafe embeddings. However, note that this transformation affects not only unsafe embeddings but also other parts of the embedding space. Therefore, an additional mechanism is required to preserve other embeddings.
3.2.2 Safe Embedding Preservation (SEP)

### v2

# Mitigating Sexual Content Generation via Embedding Distortion in Text-conditioned Diffusion Models

@@ Abstract
Diffusion models show remarkable image generation performance following text prompts, but risk generating sexual contents. Existing approaches, such as prompt filtering, concept removal, and even sexual contents mitigation methods, struggle to defend against adversarial attacks while maintaining benign image quality. In this paper, we propose a novel approach called Distorting Embedding Space (DES), a text encoder-based defense mechanism that effectively tackles these issues through innovative embedding space control. DES transforms unsafe embeddings, extracted from a text encoder using unsafe prompts, toward carefully calculated safe embedding regions to prevent unsafe contents generation, while reproducing the original safe embeddings. DES also neutralizes the “nudity” embedding, by aligning it with neutral embedding to enhance robustness against adversarial attacks. As a result, extensive experiments on explicit content mitigation and adaptive attack defense show that DES achieves state-of-the-art (SOTA) defense, with attack success rate (ASR) of 9.47% on FLUX.1, a recent popular model, and 0.52% on the widely adopted Stable Diffusion v1.5. These correspond to ASR reductions of 76.5% and 63.9% compared to previous SOTA methods, EraseAnything and AdvUnlearn, respectively. Furthermore, DES maintains benign image quality, achieving Fréchet Inception Distance and CLIP score comparable to those of the original FLUX.1 and Stable Diffusion v1.5.
@@ 1 Introduction
Recent advances in diffusion models [ddpm, diffusion], including Stable Diffusion (SD) [stablediffusion] and DALL-E [dalle3], have demonstrated remarkable capabilities in various image generation tasks such as text-to-image (T2I) synthesis and text-based image editing [instructpix2pix]. However, these models can be misused to generate deepfakes, pornographic, and Not-Safe-For-Work (NSFW) content, as highlighted by the Internet Watch Foundation [iwf]. To prevent such misuse, simple filtering-based approaches, such as blacklist-based text filtering and image-based filtering [safetychecker], can be considered as possible solutions. However, they can be readily bypassed by malicious prompts that avoid explicit keywords or leverage adversarial attacks [mma, sneaky].
Recently, several defense approaches have been proposed, such as concept removal [esd, sld] and sexual content mitigation methods [safegen], to address these vulnerabilities. However, they do not show remarkable performance in suppressing explicit content, as measured by attack success rate (ASR), or struggle to preserve benign image quality, in terms of Fréchet Inception Distance (FID) [shielddiff, safegen, unlearndiffatk], as shown in Figure LABEL:fig:teaser_asr_fid. Concept removal was often tackled by altering the U-Net, whereas more recent research tends to focus on modifying just the text encoder [safeclip, advunlearn]. One possible reason for this paradigm shift is that concept-related parameters are spread across U-Net layers [local_edit_knowledge], making it difficult to remove a concept without affecting others. In contrast, text encoder-based approaches are promising, as distinct attributes are stored in localized components [local_edit_knowledge, advunlearn].
Figure 1: Performance comparison and conceptual diagram of our approach. (a) Proposed approach offers the best performance in ASR and FID [fid], while also being cost-effective in training. The relative circle sizes indicate training time. ASRs are averaged over multiple unsafe prompts, such as Sneaky [sneaky], MMA [mma], I2P [sld], Ring-A-Bell [ringabell], and P4D [p4d]. (b) Our approach distorts the unsafe embedding space by transforming unsafe embeddings into safe regions, ensuring that embeddings derived from unsafe or adversarial prompts result in benign content generation.
@@
Based on previous studies, we also propose a method that modifies the text encoder. Furthermore, in continual learning, the opposite of unlearning, maintaining feature positions in the feature space alleviates catastrophic forgetting [lessforget]. This raises the question of whether controlling features away from their original positions could be effective for removing such unsafe information. Based on this insight, we propose a novel text encoder-based approach, Distorting Embedding Space (DES), a defense framework that satisfies both robust protection against sexual content generation and high-quality benign content generation. Unlike existing methods that struggle with implicit representations, DES uniquely controls the embedding space to capture implicit sexuality, effectively defending against explicit sexual content generation, and adversarial attacks. Our framework first transforms unsafe embeddings into a designated safe region. Since this transformation can potentially affect safe embeddings and degrade benign image generation quality, DES simultaneously trains the text encoder to reproduce the original safe embeddings, as illustrated in Figure LABEL:fig:concept. As demonstrated in Figure LABEL:fig:teaser_asr_fid, DES significantly outperforms existing defense mechanisms in terms of ASR and FID. Furthermore, DES offers remarkable efficiency in both training and inference: it requires only 90 seconds for training and introduces zero inference overhead.
@@
Our contributions are as follows: 1. We propose DES, a novel defense framework that controls the text embedding space, achieving state-of-the-art (SOTA) performance with ASRs of 9.47% on FLUX.1 and 0.52% on SDv1.5. These represent ASR reductions of 76.5% and 63.9% compared to previous SOTA, EraseAnything [eraseanything] and AdvUnlearn [advunlearn], respectively. Importantly, DES also maintains benign image generation quality. 2. We develop a practical solution that requires efficient training with zero-inference overhead, enabling easy deployment in real-world applications. 3. We conduct extensive evaluations against both explicit prompts and adversarial attacks in T2I and image-to-image (I2I) tasks. In addition, our analyses of embedding space distortion offer valuable insights into the behavior and effectiveness of DES.
@@ 2.1 Adversarial Attacks
Text-conditioned diffusion models can generate inappropriate content when given unsafe prompts [sld]. While prompt filtering can block such prompts, recent studies reveal that adversarial attacks can bypass these filters [p4d, daca, cce]. These attacks have become increasingly sophisticated, employing various optimization techniques to circumvent safety filters. SneakyPrompt [sneaky] leverages reinforcement learning to craft adversarial prompts that generate images semantically similar to target prompts, MMA-diffusion [mma] employs gradient-based optimization to create prompts that closely resemble target prompts. Ring-A-Bell [ringabell] uses a genetic algorithm to discover prompts similar to combinations of normal embeddings and extracted concept embedding. These attacks effectively bypass safety filters by exploiting unsafe embedding subspaces inherited from uncurated training data.
@@ 2.2 Defense Methods
2.2.1 Filtering-based Defense Methods
Several defense mechanisms have been proposed to address these vulnerabilities [aws_comprehend, openai_moderation], generally relying on embedding-based contextual analysis [guardt2i, safree]. GuardT2I [guardt2i] leverages a Large Language Model for NSFW detection through embedding interpretation, while SAFREE [safree] proposes training-free filtering based on distances between masked embeddings and unsafe concepts. However, these methods require additional model training or introduce inference overhead. Furthermore, these approaches struggle to detect unsafe content in ambiguous expressions and remain vulnerable to white-box attacks. In contrast, DES operates directly on the text encoder without requiring additional models or computational overhead, while effectively handling unsafe prompts in both white-box and black-box scenarios through its embedding space control.
2.2.2 Concept Removal-based Defense Methods
Recent approaches explore machine unlearning [uce, eraseanything, gloce, spm, fmn]. ESD [esd] develops a concept erasure mechanism that steers model outputs away from specific concepts. SalUn [salun] proposes saliency-based unlearning, which assigns random concepts to specific concepts to unlearn the concept. However, these UNet-based methods remain vulnerable to adversarial attacks [conceal, unlearndiffatk] or compromise image generation quality. AdvUnlearn [advunlearn] attempts to address these issues by optimizing text encoder, incorporating adversarial training. Nevertheless, it suffers from degraded image quality, a common limitation of adversarial training that compromises model performance [adv_tradeoff]. In contrast, DES overcomes these limitations through embedding space control rather than UNet modification or adversarial training, achieving robust defense while maintaining generation quality.
2.2.3 Sexual Content Mitigation Methods
SafeGen [safegen] attempts to prevent sexual content generation by fine-tuning the self-attention layers of UNet, pushing sexual content toward a blurred mosaic target using vision-only input. This text-agnostic design achieves high nudity removal rates but it introduces visible artifacts with over-censored benign contents. ShieldDiff [shielddiff] employs LoRA fine-tuning with reinforcement learning guided by a score from NudeNet [nudenet] and CLIP [clip], yet it has not been evaluated against white-box adversarial attacks. These limitations motivate our DES, which shows robustness against adaptive white-box attacks while maintaining benign image quality.
Figure 2: Overview of DES framework. During target vector generation phase, DES searches safe-unsafe vector pairs and creates target vectors by subtracting “nudity” direction from minimum similarity safe vectors. In training phase, DES aligns unsafe vectors with target vectors and maintains safe vectors by aligning both their current and nudity-integrated states with the originals. It also aligns the “nudity” vector with a neutral vector, removing its semantics. Here, $v$ and $\tilde{v}$ denote vectors from the original and training text encoders, respectively.
@@ 3 Proposed Methods
Figure 2 provides an overview of DES, illustrating the target vector generation and training phases. The first phase calculates transformation targets that redirect unsafe prompts to safe regions without disrupting safe embeddings. During the training phase, the text encoder is fine-tuned to unlearn unsafe information while preserving safe semantics.
@@ 3.1 Target Vector Generation Phase
To prevent sexual content generation, we propose transforming unsafe embeddings into the safe embedding region or to locations significantly different from their originals. This phase involves identifying optimal target safe vectors that are most dissimilar to unsafe vectors, as greater dissimilarity is assumed to enhance robustness by increasing embedding space distortion. We search through all safe vectors to identify those with minimum cosine similarity to each unsafe vector. Robustness analysis based on dissimilarity is provided in Appendix C.3. This selection procedure is formalized as:
${s}_{i}^{*}=\arg\min_{{s}_{i}}\left(\frac{{u}_{i}\cdot{s}_{i}}{\|{u}_{i}\|\|{s}_{i}\|}\right),$ (1)
where $i\!=\!1,\ldots,M$ indexes $M$ vectors, and ${s}_{i}$, ${u}_{i}$, and ${s}_{i}^{*}$ denote safe, unsafe, and selected safe vectors, respectively. Examples of selected safe prompts are provided in Appendix C.2.
We then observe the similarity between the selected safe vectors and the “nudity” vector (e.g., nudity). Interestingly, as shown in Figure 3, selected safe vectors have positive correlations
Figure 3: Cosine similarity distributions between $n$ and other vectors. Selected safe vectors initially exhibit positive similarities, which decrease as the $\frac{n}{\|n\|}$, scaled by $\alpha$, is subtracted.
with “nudity” vector. While the selection strategy improves robustness, we propose further enhancement by subtracting the “nudity” direction while using the selected vectors as basis vectors, creating target vectors $t_{i}$ that are anti-correlated with the “nudity” vector. This subtraction step is represented as:
$t_{i}={s}_{i}^{*}-\alpha\frac{n}{\|n\|},$ (2)
where $n$ is the “nudity” vector and $\alpha$ is a scaling factor. This ensures that unsafe vectors are directed away from the “nudity” direction. However, excessive subtraction can cause performance degradation if the embeddings deviate too far from the learned embedding space. Therefore, we adopt $\alpha$ to control the scale of subtraction. The green-hued distributions in Figure 3 demonstrate the successful creation of these anti-correlated vectors, which serve as more effective transformation targets for unsafe vectors. We provide a detailed description of the target vector generation phase in Algorithm 1.
@@
0: Pretrained text encoder $\mathcal{E}_{\phi_{o}}$, safe prompts $\mathcal{P}_{s}=\{p_{s,1},\ldots,p_{s,M}\}$, unsafe prompts $\mathcal{P}_{u}=\{p_{u,1},\ldots,p_{u,M}\}$, nudity prompt $p_{n}$, scale factor $\alpha$
[…]
@@
4:  $u_{i}\leftarrow\mathcal{E}_{\phi_{o}}(p_{u,i})$ // Extract unsafe vectors
5:  ${s}_{i}^{*}$ is computed by Eq. (1) // Select safe vectors
6:  $t_{i}$ is computed by Eq. (2) // Nudity subtraction
7:  $\mathcal{D}\leftarrow\mathcal{D}\cup\{(t_{i},p_{u,i},p_{s,i})\}$ // Save pairs
@@ 3.2 Training Phase
3.2.1 Distorting Unsafe Embedding Space
@@
In the text embedding space, unsafe embeddings should not occupy positions associated with unsafe content. They should be transformed into safe embedding regions or moved from their original positions by fine-tuning the text encoder. Specifically, we propose the unsafe loss, which maximizes the cosine similarity between the current unsafe vectors $\tilde{u}_{i}$ and the target safe vectors $t_{i}$:
$\mathcal{L}_{u}=\frac{1}{B}\sum_{i=1}^{B}\left(1-\frac{\tilde{u}_{i}\cdot t_{i}}{\|\tilde{u}_{i}\|\|t_{i}\|}\right),$ (3)
@@
where $i\!=\!1,\ldots,B$ represents each embedding in a mini-batch, and $B$ denotes the batch size for each iteration. Each of $\tilde{u}_{i}$ and $t_{i}$ represents an $i$-th embedding vector in a mini-batch. It aligns unsafe vectors with target vectors, avoiding their original positions. In particular, unsafe vectors become anti-correlated with the “nudity”, ensuring its removal from unsafe embeddings. However, note that this transformation affects not only unsafe embeddings but also other parts of the embedding space. Therefore, an additional mechanism is required to preserve other embeddings.
3.2.2 Safe Embedding Preservation